A 12 foot ladder is leaning against a wall. If the ladder reaches 8 feet high on the wall, what is the angle the ladder forms with the ground to the nearest degree?
Answer by tutorcecilia(2152)

Using the formula:
Sin(theta) = Opposite/Hypotenus
Sin(theta) = 8/12 [Plug-in the values]
Sin(theta) = 2/3 [Reduce the fraction]
Sin (41.81) = 2/3 [Use a calculator or a table to find sin(theta)]
41.81 degrees [Final answer]